# RaGS: Unleashing 3D Gaussian Splatting from 4D Radar and Monocular Cues  for 3D Object Detection

### v2

@@ Introduction
RaGS (Ours) 544×960 C+R 35.57 40.94
Table 3: Comparison of 3D object detection on the OmniHD-Scenes (Zheng et al. 2024) test set.
Detection Head and Loss Function
@@
On one hand, following (Bai et al. 2024), we apply the depth loss $\mathcal{L}_{\text{depth}}$ and the perspective segmentation loss $\mathcal{L}_{\text{seg}}$ to supervise the raw feature extraction for pretraining, which can be expressed as
@@
On the other hand, we adopt the 3D object detection loss $\mathcal{L}_{\text{det}}$ following (Zheng et al. 2023), along with auxiliary rendering losses: the rendered depth loss in the perspective view $\mathcal{L}_{\text{depth\_render}}$, and the semantic segmentation loss in the BEV $\mathcal{L}_{\text{seg\_render}}$. The total joint training loss is defined as
@@
where $\lambda$ is the hyperparameter balancing detection and auxiliary tasks. In this work, we simply set $\lambda=0.1$. See supplementary material for more details.
Experiments
Implementation Details
Datesets and Evaluation Metrics. We evaluate our model on three 4D radar benchmarks: the VoD (Palffy et al. 2022), TJ4DRadSet (Zheng et al. 2022), and OmniHD-Scenes (Zheng et al. 2024). The VoD dataset, collected in Delft, includes 5,139 training frames, 1,296 validation frames, and 2,247 test frames. The TJ4DRadSet, captured in Suzhou, contains 7,757 frames with 5,717 for training and 2,040 for testing, posing challenges due to complex scenarios such as nighttime and glare. The OmniHD-Scenes dataset features multimodal data from six cameras, six 4D radars, and a 128-beam LiDAR, with 11,921 keyframes annotated, consisting of 8,321 for training and 3,600 for testing.
@@
Network Settings and Training details. For all datasets, the final voxel is set to a cube of size 0.32 m, and the image sizes are 800×1280 for VoD, 640×800 for TJ4DRadSet, and 544×960 for OmniHD-Scenes. Anchor size and point cloud range are kept as in (Zheng et al. 2024). The models are trained on 4 NVIDIA GeForce RTX 4090 GPUs with a batch size of 4 per GPU. AdamW is used as the optimizer, with 12 epochs for pretraining and 24 epochs for joint training.
3D Object Detection Results
Results on View-of-Delft and TJ4DRadSet Dataset. Table 1 presents the 3D object detection results on the VoD validation set (Palffy et al. 2022), where our RaGS consistently outperforms state-of-the-art 4D radar and camera fusion models. For cars, we achieve exceptional accuracy, leveraging the flexible occupancy capability of the Gaussian model and an effective fusion design. For pedestrians and cyclists, we also deliver near best performance. When compared to the strong BEV-based baseline LXL (Xiong et al. 2024), our method shows significant improvements, with gains of 5.55% in $\text{mAP}_{\text{EAA}}$ and 6.57% in $\text{mAP}_{\text{DC}}$. Furthermore, it achieves a frame rate of 10.5 FPS, demonstrating its suitability for real-world applications. Compared to the VoD dataset, the TJ4DRadSet dataset (Zheng et al. 2022) presents greater challenges, including complex scenarios such as nighttime, under-bridge conditions, and the addition of a truck category with highly variable object sizes. Despite these difficulties, RaGS achieves state-of-the-art performance, with the highest $\text{mAP}_{\text{BEV}}$ of 51.04% and the second-best $\text{mAP}_{\text{3D}}$ of 41.72%, as shown in Table 2. Visualization results of VoD and TJ4DRadSet are shown in the first and second row of Fig. 5, respectively. Qualitative comparison are presented in Fig. 6.
@@
Results on OmniHD-Scenes Dataset. We also evaluated the performance of RaGS in a surround-view scenario, as shown in Table 3. While RaGS exhibits a slight performance drop in terms of the ODS metric, it consistently outperforms traditional fixed-grid models RCfusion and BEVFusion on the mAP metric. This underscores the advantage of utilizing Gaussian representations for the fusion of 4D radar and surround-view data in the context of 3D object detection.
Ablation Study
Overall Ablation. Table 4 presents an ablation study evaluating the contributions of FLI, IMA, and MGF, using Gaussianformer (Huang et al. 2024) with 4D radar as our baseline. FLI significantly aids Gaussian initialization through depth back-projection and foreground pixel selection. IMA effectively combines multimodal information, as shown in Fig. 4, progressively updating Gaussian attributes to reconstruct sparse regions of interest efficiently. Finally, MGF enhances performance by fusing multi-level features from Gaussian rasterization, enabling multi-layered perception and improving overall results.
Ablation on FLI. As shown in Table 5, FLI consists of three components: frustum-based initialization, integration of 4D radar, and depth back-projection. The frustum-based initialization helps avoid resource wastage from out-of-view objects, while the 4D radar and depth back-projection provide prior knowledge, reducing the impact of unstable segmentation and depth estimation. Additionally, the coarse but structured initialization from FLI is further refined through IMA for eliminating instability.
@@
Ablation on Hyper-parameters and Losses. We conduct an ablation study on the fusion level $L$ in MGF, the aggregation number $M$ in IMA, and the anchor number $N$. Both mAP and GFLOPS are reported. To balance performance and computational cost, we select 2-level MGF fusion, 3 layers for IMA, and 12,800 Gaussian anchors. See Table S1 and S2 in the supplementary material for further details.
@@
RCFusion 49.65 33.85 55.3M 9.0 680.8
LXL 56.31 36.32 64.0M 6.1 1169.2
SGDet3D 59.75 41.82 73.6M 9.2 1309.3
RaGS 61.86 41.72 56.5M 10.5 781.0
@@
Comparison with State-of-the-art Models. We compared RaGS with existing 4D radar and camera fusion models on mAP, FPS, and GFLOPS. As shown in Table 7, RaGS outperforms with fewer parameters and lower GFLOPS. Unlike BEV-grid methods, RaGS uses iterative refinement to focus on sparse objects while maintaining scene understanding. SGDet3D uses 160$\times$160$\times$8 fixed sampling points, about 16 times more than our 12,800. Table 8 presents efficiency analysis of each component within the RaGS Framework. The Gaussian in IMA is resource-intensive due to dynamic object focus, and future work will improve efficiency.
Conclusions
In this work, we propose RaGS, the first framework to leverage 3D Gaussian Splatting for fusing 4D radar and monocular images in 3D object detection. RaGS models the fused multi-modal features as continuous 3D Gaussians, enabling dynamically object attention while preserving comprehensive scene perception. The cascaded architecture, consisting of Frustum-based Localization Initiation (FLI), Iterative Multimodal Aggregation (IMA), and Multi-level Gaussian Fusion (MGF), progressively constructs and refines the Gaussian field through frustum-based localization, cross-modal aggregation, and multi-level fusion. Experiments on public benchmarks show state-of-the-art performance.
Limitations. RaGS currently works on single-frame data. Ongoing research aims to incorporate sequential modeling and utilize 4D radar for more explicit object attention.

### v3

@@ 3.2 Gaussian Aggregation
RaGS (Ours) 544×960 C+R 35.88 43.45
Table 1: Comparison on the OmniHD-Scenes [OmniHD] test set.
Approach Input $\text{AP}_{\text{3D}}$ (%) $\text{AP}_{\text{BEV}}$ (%)
ImVoxelNet (WACV 2022) [imvoxelnet] C 14.96 17.12
RadarPillarNet (TIM 2023) [RCFusion] R 30.37 39.24
SMURF (TIV 2023) [SMURF] R 32.99 40.98
FUTR3D (CVPR 2023) [FUTR3D] R+C 32.42 37.51
BEVFusion (ICRA 2023) [BEVFusion] R+C 32.71 41.12
LXL (TIV 2024) [LXL] R+C 36.32 41.20
RCFusion (TIM 2023) [RCFusion] R+C 33.85 39.76
LXLv2 (RAL 2025) [LXLv2] R+C 37.32 42.35
UniBEVFusion (ICRA 2025) [UniBEVFusion] R+C 37.76 42.92
MSSF (TITS 2025) [MSSF] R+C 37.97 43.11
HGSFusion (AAAI 2025) [HGSFusion] R+C 37.21 43.23
SGDet3D (RAL 2025) [SGDet3D] R+C 41.82 47.16
RaGS (Ours) R+C 41.95 51.04
Table 2: Comparison on the test set of TJ4DRadSet [TJ4D].
Models View-of-Delft[VoD] TJ4DRadSet[TJ4D]
$\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$ $\text{mAP}_{\text{3D}}$ $\text{mAP}_{\text{BEV}}$
FUTR3D (CVPR2023) [FUTR3D] 49.03 69.32 32.42 37.51
RaCFormer (CVPR2025) [racformer] 54.44 78.57 - -
RaGS (Ours) 61.86 81.63 41.95 51.04
Table 3: Comparison with DETR-based models.
Approach Input Entire Annotated Area (%) Driving Corridor (%) FPS
Car Ped Cyclist mAP Car Ped Cyclist mAP
ImVoxelNet (WACV 2022) [imvoxelnet] C 19.35 5.62 17.53 14.17 49.52 9.68 28.97 29.39 11.1
PillarNeXt (CVPR 2023) [pillarnext] R 30.81 33.11 62.78 42.23 66.72 39.03 85.08 63.61 -
RadarPillarNet (TIM 2023) [RCFusion] R 39.30 35.10 63.63 46.01 71.65 42.80 83.14 65.86 98.8
CenterPoint (CVPR 2021) [centerpoint] R 35.84 41.03 67.11 47.99 70.65 50.14 85.67 68.82 38.3
VoxelNeXt (CVPR 2023) [VoxelNeXt] R 36.98 42.37 68.15 49.17 70.95 51.85 87.33 70.04 31.6
SMURF (TIV 2023) [SMURF] R 42.31 39.09 71.50 50.97 71.74 50.54 86.87 69.72 -
SCKD (AAAI 2025) [SCKD] R 41.89 43.51 70.83 52.08 77.54 51.06 86.89 71.80 39.3
FUTR3D (CVPR 2023) [FUTR3D] R+C 46.01 35.11 65.98 49.03 76.98 43.10 86.19 69.32 7.3
BEVFusion (ICRA 2023) [BEVFusion] R+C 37.85 40.96 68.95 49.25 70.21 45.86 89.48 68.52 7.1
RCFusion (TIM 2023) [RCFusion] R+C 41.70 38.95 68.31 49.65 71.87 47.50 88.33 69.23 9.0
RCBEVDet (CVPR 2024) [RCBEVDet] R+C 40.63 38.86 70.48 49.99 72.48 49.89 87.01 69.80 -
ZFusion (CVPR 2025) [zfusion] R+C 43.89 39.48 70.46 51.28 79.51 52.95 86.37 72.94 -
IS-Fusion (CVPR 2024) [IS-Fusion] R+C 48.57 46.17 68.48 54.40 80.42 55.50 88.33 74.75 -
LXL (TIV 2024) [LXL] R+C 42.33 49.48 77.12 56.31 72.18 58.30 88.31 72.93 6.1
SGDet3D (RAL 2025) [SGDet3D] R+C 53.16 49.98 76.11 59.75 81.13 60.91 90.22 77.42 9.2
RaGS (Ours) R+C 58.15 50.81 76.62 61.86 88.15 61.67 95.07 81.63 10.5
Table 4: Comparison with state-of-the-art approaches on the val set of View-of-Delft [VoD].
Multi-level Gaussian Fusion (MGF). After obtaining multiple Gaussian representations that capture different levels of semantic and geometric abstraction, we further integrate them to form a unified, multi-level feature representation for subsequent rendering and detection. Through IMA, we obtain Gaussians set consisting of $M$ refined Gaussians $\mathcal{G}=\{\mathbf{G}_{i}\}_{i=1}^{M}$. Each $\mathbf{G}_{i}$ has $N$ Gaussian, inheriting its explicit attributes $(\mathbf{P}_{i},\mathbf{R}_{i},\mathbf{S}_{i},\mathbf{O}_{i})$ and implicit feature $\mathbf{F}^{\text{I}}_{i}$ from the previous stage, which can be formally represented as $\mathbf{G}_{i}=(\bm{\mu}_{i},\bm{\Sigma}_{i},\alpha_{i},\mathbf{f}_{i})$, where the Gaussian center $\bm{\mu}_{i}$ corresponds to $\mathbf{P}_{i}$, the covariance $\bm{\Sigma}_{i}$ is derived from the rotation $\mathbf{R}_{i}$ and scale $\mathbf{S}_{i}$, $\alpha_{i}$ denotes opacity (from $\mathbf{O}_{i}$), and $\mathbf{f}_{i}$ is the learned feature embedding.
For each of $N$ Gaussian within $\mathbf{G}_{i}$, it contributes to BEV cells through differentiable Gaussian Splatting. Specifically, given a BEV pixel at position $\mathbf{q}=(x,y)$, its feature value is computed as the weighted accumulation of all projected Gaussians, formulated as
$\sum_{n=1}^{N}\alpha_{n}\exp\!\left(-\tfrac{1}{2}(\mathbf{q}-\bm{\mu}_{n,xy})^{\!\top}\!\bm{\Sigma}_{n,xy}^{-1}(\mathbf{q}-\bm{\mu}_{n,xy})\right)\mathbf{f}_{n},$ (9)
where $\bm{\mu}_{n,xy}$ and $\bm{\Sigma}_{n,xy}$ are the mean and covariance of the Gaussian projected onto the BEV plane. This operation defines a smooth, differentiable mapping from the 3D Gaussian field to the 2D BEV domain, implemented efficiently via a CUDA-based rasterizer [Gaussianformer]. The implicit features of the last $L~(L\!\leq\!M)$ Gaussians are rasterized into multi-level BEV-aligned feature maps $\{\mathbf{F}^{(l)}\in\mathbb{R}^{X\times Y\times C}\}$ as
$\mathbf{F}^{(l)}=\mathtt{Rasterize}(\mathbf{G}_{M-L+l}),\quad l=1,\ldots,L,$ (10)
where $(X,Y)$ denotes the BEV resolution, $\mathtt{Rasterize}$ denotes the rasterizer. These rendered maps form a hierarchical Gaussian representation across spatial scales. We then fuse the multi-level features via convolution as
$\mathbf{F}^{\text{gs}}=\mathtt{Conv}\!\left(\mathtt{Concat}(\mathbf{F}^{(1)},\ldots,\mathbf{F}^{(L)})\right).$ (11)
The fused Gaussian feature $\mathbf{F}^{\text{gs}}$ is further enhanced by radar-derived sparse pillars $\mathbf{F}^{\text{pillar}}$ through a cross-modal fuser, formulated as $\mathbf{F}^{\text{BEV}}=\mathtt{CMF}(\mathbf{F}^{\text{gs}},\mathbf{F}^{\text{pillar}})$, where $\mathtt{CMF}$ denotes a lightweight module composed of concatenation and convolution layers. As a result, our MGF generates a hierarchical representation that compensates for the sparse and low-resolution nature of the Gaussian from IMA.
Additionally, we render final-layer Gaussians into depth supervised by LiDAR, and predict a BEV segmentation map from fused BEV features with occupancy guidance. Crucially, while MGF rasterizes the Gaussian field into BEV space, the Gaussians themselves are dynamically refined toward object-centric regions, which allows adaptive aggregation that prioritizes sparse foreground objects and mitigates the background redundancy of grid-based BEV methods.
3.3 Detection Head and Loss Function
@@
On one hand, following [SGDet3D], we apply the depth loss $\mathcal{L}_{\text{depth}}$ and the perspective segmentation loss $\mathcal{L}_{\text{seg}}$ to supervise the raw feature extraction for pretraining, expressed as
@@
On the other hand, we adopt the 3D object detection loss $\mathcal{L}_{\text{det}}$ following [RCFusion], along with auxiliary rendering losses: the rendered depth loss in the perspective view $\mathcal{L}_{\text{depth\_render}}$, and the semantic segmentation loss in the BEV $\mathcal{L}_{\text{seg\_render}}$. The total joint training loss is defined as
@@
where $\lambda$ is the hyperparameter balancing detection and auxiliary tasks. In this work, we simply set $\lambda=0.1$. The perspective and BEV segmentation ground-truths can be simply generated through detectron2 [wu2019detectron2] and 3D bounding boxes. See supplementary material for more details.
3.4 Implementation Details
Datesets and Evaluation Metrics. We evaluate our model on three 4D radar benchmarks: VoD [VoD], TJ4DRadSet [TJ4D], and OmniHD-Scenes [OmniHD]. The VoD dataset, collected in Delft, includes 5,139 training frames, 1,296 validation frames, and 2,247 test frames. The TJ4DRadSet, captured in Suzhou, contains 7,757 frames with 5,717 for training and 2,040 for testing, posing challenges due to complex scenarios such as nighttime and glare. The OmniHD-Scenes dataset features multimodal data from six cameras, six 4D radars, and a 128-beam LiDAR, with 11,921 keyframes annotated, consisting of 8,321 for training, 3,600 for testing.
@@ 3.4 Implementation Details
Network Settings and Training details. For all datasets, the final voxel is set to a cube of size 0.32 m, and the image sizes are 800×1280 for VoD, 640×800 for TJ4DRadSet, and 544×960 for OmniHD-Scenes. Anchor size and point cloud range are kept as in [OmniHD]. The models are trained on 4 NVIDIA GeForce RTX 4090 GPUs with a batch size of 4 per GPU. AdamW is used as the optimizer, with 12 epochs for pretraining and 24 epochs for joint training.
@@
4 Experiments
4.1 3D Object Detection Results
Results on OmniHD-Scenes Dataset. We evaluated the performance of RaGS in a surround-view scenario, as shown in Table 1. RaGS consistently outperforms traditional fixed-grid models RCFusion and BEVFusion on the mAP and ODS. This underscores the advantage of utilizing Gaussian representations for the fusion of 4D radar and surround-view data in the context of 3D object detection. For multi-view setting, our ongoing work is exploring more elegant initialization prior like VGGT [VGGT].
@@ 4 Experiments
NN Gaussian Anchor $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$ FLOPS
Car Ped Cyc
3200 52.24 41.79 70.29 54.77 76.13 599.39
6400 52.23 41.92 76.35 56.83 76.22 611.95
12800 52.94 48.31 77.10 59.45 76.98 639.82
19200 52.63 48.84 76.94 59.47 76.70 666.06
Table 8: Ablation of anchor number NN.
Results on TJ4DRadSet and View-of-Delft Dataset. Compared to the VoD dataset, the TJ4DRadSet dataset [TJ4D] presents greater challenges, including complex scenarios such as nighttime, under-bridge conditions, and the addition of a truck category with highly variable object sizes. Despite these difficulties, RaGS achieves state-of-the-art performance, with the highest $\text{mAP}_{\text{BEV}}$ of 51.04% and the best $\text{mAP}_{\text{3D}}$ of 41.95%, as shown in Table 2. On the other side, Table 4 presents the 3D object detection results on the VoD validation set [VoD], where our RaGS consistently outperforms state-of-the-art 4D radar and camera fusion models. For cars, we achieve exceptional accuracy, leveraging the flexible occupancy capability of the Gaussian model and an effective fusion design. For pedestrians and cyclists, we also deliver near best performance. When compared to the strong BEV-based baseline LXL [LXL], our method shows significant improvements, with gains of 5.55% in $\text{mAP}_{\text{EAA}}$ and 8.70% in $\text{mAP}_{\text{DC}}$. Furthermore, it achieves a frame rate of 10.5 FPS, demonstrating its suitability for real-world applications. Visualization results of VoD and TJ4DRadSet are shown in the first and second row of Fig. 5, respectively. Qualitative comparison are presented in Fig. 6. Compared to DETR-based methods such as FUTR3D (49.03%) and RaCFormer (54.44%), as shown in Table 3, our RaGS (61.86%) demonstrates a significant performance advantage. Unlike implicit end-to-end query-based frameworks, the 3D Gaussian Splatting in RaGS offers explicit physical interpretability, serving as a multi-modal aggregator that fuses semantic and geometric features, leveraging radar for 3D perception and monocular cues for enriched observation.
4.2 Ablation Study
Overall Ablation. Table 5 reports the overall ablation on RaGS, conducted with half training epochs for efficiency. Starting from the baseline [Gaussianformer] with 4D radar input, each module consistently contributes to performance improvement. FLI provides reliable Gaussian initialization via depth-guided localization, IMA progressively refines multimodal aggregation, and MGF complements them with multi-level feature fusion. Together, these modules form a coherent framework that balances accuracy and efficiency while ensuring stable performance across different conditions.
Ablation on FLI. As shown in Table 6, FLI consists of three components: frustum-based initialization, integration of 4D radar, and depth back-projection. The initialization helps avoid resource wastage from out-of-view objects, while the 4D radar and depth back-projection provide prior knowledge, reducing the potential impact of unstable segmentation and depth estimation. Additionally, the coarse but structured Gaussian from FLI will be refined through IMA for eliminating instability.
@@ 4.2 Ablation Study
Disturbance $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$
LXL RaGS LXL RaGS
$\pm$ 0 , $\pm$ 0.0m 56.42 61.86 73.03 81.63
$\pm$ 2 , $\pm$ 0.2m 56.26 61.75 73.16 81.42
$\pm$ 5 , $\pm$ 0.5m 50.25 56.66 72.07 76.39
Table 9: Performance comparison under calibration disturbance.
Condition $\text{mAP}_{\text{EAA}}$ $\text{mAP}_{\text{DC}}$
LXL RaGS LXL RaGS
Rain 51.03 56.79 72.62 75.96
Haze 51.16 56.72 69.49 72.29
Lowlight 52.13 56.89 70.26 76.02
Table 10: RaGS performance under simulated adverse weather.
Figure 6: Qualitative comparison with state-of-the-art.
@@
RCFusion [RCFusion] 49.65 33.85 55.3M 9.0 680.8
LXL [LXL] 56.31 36.32 64.0M 6.1 1169.2
HGSFusion [HGSFusion] 58.96 37.21 64.6M 3.2 1899.6
SGDet3D [SGDet3D] 59.75 41.82 73.6M 9.2 1309.3
RaGS (ours) 61.86 41.95 56.5M 10.5 781.0
@@
Ablation on Hyper-parameters. We conduct an ablation study on the the anchor number $N$, as shown in Table 8, where both mAP and GFLOPS are reported. To balance performance and computational cost, we select 2-level MGF fusion, 3 layers for IMA, and 12,800 Gaussian anchors. For fusion level $L$ in MGF and the aggregation number $M$ in IMA, see supplementary material for further hyperparameter ablation details.
4.3 Robustness Analysis
[…]
4.4 Computational Cost
We compared RaGS with existing 4D radar and camera fusion models on mAP, FPS, and GFLOPs. As shown in Table 11, RaGS outperforms with fewer parameters and lower GFLOPS. Unlike BEV-grid methods, RaGS uses iterative refinement to focus on sparse objects while maintaining scene understanding. SGDet3D uses $\times160$$\times$8 fixed sampling points, about 16 times more than our 12,800. Table 12 presents efficiency analysis of each component within the RaGS Framework.
5 Conclusions
In this work, we propose RaGS, the first framework to leverage 3D Gaussian Splatting for fusing 4D radar and monocular images in 3D object detection. RaGS models the fused multi-modal features as continuous 3D Gaussians, enabling dynamically object attention while preserving comprehensive scene perception. The architecture, consisting of FLI, IMA, and MGF, progressively constructs and refines the Gaussian field through frustum-based localization, cross-modal aggregation, and multi-level fusion. Experiments on public benchmarks show state-of-the-art performance.
Limitations. Our ongoing work focuses on extending it to sequential modeling by explicitly exploiting radar velocity within 4D GS for enhanced temporal object awareness.